### v1

# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

Abstract.
Material defects represent a primary challenge affecting product performance and giving rise to safety issues in related products. The rapid and accurate identification and localization of material defects constitute crucial research endeavours in addressing contemporary challenges associated with material defects. Although conventional non-destructive testing methods such as ultrasonic and X-ray approaches have mitigated issues related to low efficiency in manual inspections, they struggle to meet the diverse requirements of high precision, real-time speed, automation, and intelligence. In recent years, propelled by the swift advancement of machine learning technologies, particularly exemplified by deep learning, machine learning has swiftly emerged as the core technology and a prominent research direction for material defect detection. Through a comprehensive review of the latest literature, we systematically survey the machine learning techniques applied in material defect detection into five categories: unsupervised learning, supervised learning, semi-supervised learning, reinforcement learning, and generative learning. We provide a detailed analysis of the main principles and techniques used, together with the advantages and potential challenges associated with these techniques. Furthermore, the survey focuses on the techniques for defect detection in composite materials, which are important types of materials enjoying increasingly wide application in various industries such as aerospace, automotive, construction, and renewable energy. Finally, the survey explores potential future directions in material defect detection utilizing machine learning technologies. This comprehensive survey not only consolidates existing literature on machine learning-based material defect detection technologies but also serves as a foundational reference for future researchers and industrial practitioners, providing valuable insights and guidance in developing advanced and efficient material defect detection systems.
Material Defect Detection, Composites Manufacturing, Machine Learning, Deep Learning, Computer Vision, Machine Vision.
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456conference: ACM Computing Surveys; ACM Computing Surveysbooktitle: ;price: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Material defect detection (MDD) stands as a critical subfield within the broader domain of defect detection, with a primary focus on identifying and assessing defects in various materials. MDD is a pivotal element in ensuring product quality and production efficiency, and it holds particular significance in the manufacturing industry. The defects can occur on the surface or internally, potentially compromising the product’s performance and safety (usamentiaga2022automated, ). As such, MDD is an essential component in numerous industries, playing a vital role in ensuring the safety, reliability, and quality assurance of products. Adherence to strict regulatory standards is crucial, not only to prevent catastrophic failures and reduce waste but also to maintain cost efficiency. Effective defect detection methodologies directly contribute to the longevity and performance of products, thereby driving innovation in manufacturing processes. Additionally, they play a key role in promoting environmental sustainability by minimizing material waste. In this regard, MDD is fundamental to the responsible and successful production and distribution of goods across various sectors.
Current MDD methods must adapt to the increasing complexity of manufacturing processes and product diversity to meet the demands of modern manufacturing. As a result, manufacturers face a series of challenges and higher detection requirements. One major challenge is the identification of complex defects. The growing variety of materials and the complexity of manufacturing processes lead to increasingly intricate defect types within materials, such as cracks, pores, and inclusions. These defects exhibit diverse shapes, sizes, and positions, making their identification difficult. Another challenge is the detection of minute defects. Some material defects are exceedingly small, even imperceptible to the naked eye, and require high-precision detection methods to identify their potentially significant impacts on product performance and safety. Furthermore, MDD often involves handling multimodal data, including images, ultrasonic waves, and X-rays. Each data type has distinct features and noise levels, posing a challenge in effectively processing and integrating these diverse sources of information. Real-time online detection of defects in materials is also crucial for timely issue identification and resolution on production lines (shi2020overview, ; ou2021recent, ). This requirement necessitates the development of methods with fast, stable, and reliable performance. The ability to detect defects in real-time enables manufacturers to promptly address issues, minimize production interruptions, and maintain high product quality standards (yang2020using, ; zeng2022small, ; zhao2023multi, ). Moreover, the rarity of defective samples also poses difficulties in collecting sufficient data for analysis. Therefore, developing effective and efficient MDD techniques that can operate optimally with these practical challenges is becoming increasingly crucial (jin2022survey, ; shahrabadi2022defect, ).
Traditional approaches to MDD often depend on manual inspection, which comes with significant limitations. As highlighted by Bhatt et al. (bhatt2021image, ), manual inspection is subject to the subjective judgment of inspectors and generally lacks precision, particularly when dealing with minuscule defects or in environments that are potentially harmful to humans. Moreover, manual inspection struggles to keep pace with the increasing complexity of products and the accelerating speeds of modern manufacturing processes. Consequently, manual inspection is often insufficient to meet the rigorous demands of contemporary production standards, necessitating the adoption of more advanced and reliable methods for surface defect detection in industrial products.
As a response to these challenges, the implementation of visual perception-based surface defect detection has gained prominence as a more effective alternative, progressively replacing manual inspections. Machine vision and computer vision, as rapidly evolving branches in the field of artificial intelligence, have found extensive applications across various industrial domains, including the critical area of MDD (hafiz2021attention, )(ren2022state, ). This advanced technology, which employs non-contact automatic detection, offers higher accuracy and demonstrates reliable performance in complex production environments. The integration of these technologies into MDD involves the automated inspection and analysis of product surfaces or interiors, utilizing advanced image processing and computer vision algorithms. These cutting-edge technologies enable the automatic identification and classification of defects, proving capable of handling vast volumes of data and conducting high-precision inspections rapidly. Moreover, the application of machine vision and computer vision significantly enhances the accuracy and consistency of defect detection. By reducing the reliance on human inspection, these technologies diminish the impact of human error, contributing to more reliable and uniform inspection results. This improvement in detection quality, in turn, helps to lower production costs and boost overall production efficiency (ren2022state, ; luo2020automated, ; fan2020line, ). The effectiveness of visual perception-based surface defect detection has been proven across various sectors. For instance, Stojanovic et al. (stojanovic2001real, ) showcased its application in the tile industry, while Wang et al. (wang2021defect, ) and Dong et al. (dong2021automatic, ) have made notable innovations in the fabric sector. Similarly, significant contributions to steel plate defect detection have been made by Hu et al. (hu2020unsupervised, ) and Wang et al. (wang2018distributed, ). Furthermore, the technology has been successfully applied in PCB defect detection, as evidenced by the works of Zhang et al. (zhang2021semi, ) and Mei et al. (mei2018unsupervised, ).
Machine learning (ML), a subset of artificial intelligence, has been the driving force behind automated defect detection based on machine vision and computer vision. Since its inception in the mid-20th century, ML has evolved significantly, from early concepts like the Turing Test and the development of learning algorithms to the adoption of statistical methods and the breakthroughs in deep learning (DL) and neural networks (janiesch2021machine, ). The advent of big data and increased computational power has led to a resurgence in the field, with applications spanning various sectors, including healthcare, finance, and autonomous vehicles (cioffi2020artificial, ). ML plays a pivotal role in MDD by automating the entire detection process, from data preprocessing and feature extraction to defect identification and classification. This automation reduces manual intervention, increases efficiency, and lowers detection costs (nturambirwe2020machine, ). Trained ML algorithms exhibit heightened sensitivity, enabling the precise identification of extremely small defects, even those imperceptible to the human eye. This significantly improves the accuracy of MDD and facilitates early issue detection, preventing declines in product quality. Moreover, ML models ensure sustained and stable detection capabilities over extended periods, unaffected by factors like fatigue or emotions, thereby enhancing reliability and consistency. These models excel at processing large volumes of data in short timeframes, efficiently handling defect detection tasks for numerous products or images and meeting the demands of large-scale production. In addition to real-time defect detection, ML technologies provide valuable data support for quality control through offline data analysis. This analysis contributes to process improvements, enhanced product quality, and the discovery of new defect patterns and trends in historical data. By uncovering these insights, ML guides scientific refinements in manufacturing processes, enabling continuous optimization and reducing the occurrence of defects (sresakoolchai2022railway, ; cao2020review, ). The integration of ML in MDD not only streamlines the detection process but also fosters a data-driven approach to quality control, ultimately leading to higher-quality products and more efficient manufacturing processes.
Despite the high detection efficiency and accuracy of ML in MDD, there are several drawbacks to consider. One significant limitation is the poor generality of detection algorithms, as they may not be readily applicable to other materials or contexts, hindering widespread adoption across diverse manufacturing settings. Another drawback is the high requirements for the quality of defect visual images, as complex backgrounds and excessive noise can increase the complexity of detection algorithms and reduce their accuracy. Obtaining high-quality, clean, and well-prepared visual data can be time-consuming and resource-intensive, especially in industrial environments with varying conditions. Furthermore, the maintenance and production costs for machine vision detection systems can be substantial due to the need for specialized hardware, software, and expertise. The initial investment and ongoing costs for upgrades, repairs, and personnel training may be a barrier for some manufacturers, particularly small and medium-sized enterprises with limited resources (cui2021sddnet, ; caiazzo2022towards, ).
With the development of new ML technologies, particularly those centred around DL, there has been an increasing application in MDD. DL, a branch of ML that utilizes artificial neural networks to simulate the operation of the human brain’s neural network, aims to process and analyze complex data (lecun2015deep, ). Common types of deep networks include: (1) Convolutional Neural Networks (CNN), which are used for processing image data, extracting features layer by layer, and demonstrating robust feature learning and classification capabilities, with common models including LeNet, AlexNet, VGGNet, GoogLeNet, and ResNet; (2) Recurrent Neural Networks (RNN), which are applied for handling sequential data such as text, speech, and time series, and can be used for real-time monitoring and prediction of continuous data on production lines in MDD; (3) Long Short-Term Memory Networks (LSTM), a special type of RNN effective in handling long sequential data, addressing the gradient vanishing problem seen in traditional RNNs, and applicable for real-time monitoring and prediction of continuous image or production line data in MDD; and (4) Generative Adversarial Networks (GAN), which comprise a generator and a discriminator that competitively achieve unsupervised learning of data, and can be employed to generate realistic defect samples, enhancing the model’s generalization capability in MDD (shrestha2019review, ).
By constructing suitable network structures, DL can effectively address the issue in traditional machine vision detection algorithms, where different algorithms are required for different MDD tasks. Various DL technologies can extract high-level features from images without the need for manual feature design, facilitating defect detection and recognition based on these features and significantly advancing the field of MDD (hafiz2021attention, ). These technologies can meet the multiple requirements of contemporary MDD, including automation, intelligence, high detection efficiency, high detection accuracy, and real-time detection (bhatt2021image, ). The proliferation of high-performance computers and GPUs has also made the application of DL in MDD more convenient and cost-effective, as DL requires substantial computational resources.
The existing literature on defect detection predominantly focuses on general methodologies or specific application domains (P2, ; P6, ; P7, ; P8, ; P10, ; P11, ), with some studies delving into defect detection in particular types of materials (P3, ; P5, ; P9, ; P17, ). However, a systematic review of ML-driven MDD remains largely unexplored. Our survey presents several distinct advantages that highlight its contributions to the field:
  1. (1)
    Firstly, we systematically gather the most recent and relevant literature regarding ML-driven MDD, thereby ensuring a comprehensive and methodical coverage of the subject matter. Our survey encompasses a broad spectrum of research works, providing a holistic view of the field;
  2. (2)
    Furthermore, we classified the various ML technologies currently employed in MDD applications. This classification is followed by a thorough analysis of each technology category. Such a comprehensive evaluation aids in discerning the suitability and effectiveness of different technological approaches in various contexts;
  3. (3)
    Moreover, our survey places special emphasis on MDD techniques specifically designed for composite materials. Composite materials are gaining increasing importance across various industries due to their unique properties and performance advantages. By dedicating attention to the challenges posed by the complex structures and heterogeneous nature of composite materials, we provide valuable insights into the state-of-the-art techniques tailored to address these specific issues;
  4. (4)
    Building on the analysis of existing technological works, we identified key challenges in current MDD practices and highlighted future directions of research in MDD. By providing these insights, we aim to guide researchers and practitioners toward addressing the limitations of current approaches and advancing the field of ML-driven MDD.
The rest of this survey paper is organized as follows. In Section 2, we present the preliminaries and the problem formulation for MDD. Following this, in Section 3, we provide a detailed technical classification of ML techniques applied in MDD. We systematically analyze and compare the technical characteristics, application domains, and pros and cons of each technology category. Section 4 focuses on the ML techniques specifically designed for defect detection in composite materials. In Section 5, we outline the open issues in current MDD practices and highlight potential research directions for the future. Finally, we conclude our paper in Section 6, summarizing the key contributions of this comprehensive survey.
2. Preliminaries and Problem Identification
In this section, we first introduce the fundamental concepts and procedural aspects of material defect detection. We provide an overview of the various types of material defects and their impact on the integrity and performance of materials. We then delve into the basic detection techniques used to identify these defects, highlighting their significance in ensuring product quality and safety. Following this, we discuss the critical role that machine learning plays in enhancing defect detection techniques, offering improvements in accuracy, efficiency, and scalability. Finally, we identify and analyze the specific challenges that current material defect detection methods face, emphasizing the need for advanced machine learning solutions to address these issues effectively.
Table 1. Common Types of Material Defects
Defect Type Description
Cracks (golewski2021validation, ) Linear separations or fractures in the material, often due to stress or impact.
Porosity (wang2022role, ) Presence of voids or pores within the material, often resulting from trapped gases during processing.
Inclusions (gong2020deep, ) Foreign materials embedded within the host material, which can compromise strength and integrity.
Corrosion (li2020point, ) Chemical or electrochemical degradation of materials, often leading to surface deterioration or structural weakness.
Delamination (segers2020backside, ) Separation of layers in laminated materials, often due to weak adhesive bonds or external stresses.
Discolouration (song2023x, ) Unintended changes in material colour, often indicative of chemical changes or degradation.
Warping (saluja2020closed, ) Deformation of the material from its intended shape, often due to uneven heat or stress distributions.
Wear (zhang2022study, ) Deterioration or degradation of materials’ physical properties over time, often as a result of repeated use, friction, or exposure to environmental factors.
2.1. Types of Material Defects
The integrity and performance of materials are pivotal in determining the reliability and longevity of engineered products and structures. A critical aspect of material science and engineering is the identification, understanding, and mitigation of material defects, which can significantly impact product quality and safety (arandiyan2021defect, ; choudhary2022recent, ). Material defects, often microscopic in nature, can escalate into macroscopic failures, leading to catastrophic outcomes in high-stakes applications such as aerospace, automotive, and civil infrastructure. The different types of material defects are listed and described in Table 1. Understanding these defects is crucial for developing strategies to detect, mitigate, or prevent them, thereby enhancing material performance and extending the lifespan of products and structures.
2.2. Defect Detection
In the manufacturing process of industrial products, various factors such as product design, processing environment, and production equipment failures often lead to the occurrence of diverse defects in the produced items, such as depressions, scratches, and voids. These defects contribute to a decline in the yield of high-quality products, escalate production costs for enterprises, and result in wastage of resources and economic losses. Importantly, defective products, once introduced into the market, pose a potential risk of causing significant safety incidents (kumbhar2023deepinspect, ; selot2023machine, ). Consequently, defect detection holds crucial value and significance in enhancing product quality and improving production efficiency, garnering joint attention from both the industrial and academic communities.
Defect detection technology primarily focuses on inspecting surface defects of test samples, such as spots, pits, scratches, colour variations, and deficiencies, and obtains relevant information about the defects, including their location, type, depth, size, contours, and other characteristic features (ren2022state, ). Previous defect detection heavily relied on manual methods, exhibiting drawbacks such as high costs, low efficiency, suboptimal accuracy, and the inability to conduct large-scale inspections. However, with the maturity of information perception sensor technologies, such as visual imaging techniques, the advancement of defect detection technologies based on machine learning has been propelled (prunella2023deep, ; usamentiaga2022automated, ). This progress has significantly addressed the requirements of modern defect detection for high precision, efficiency, complexity, scalability, and real-time capabilities, thereby holding considerable research value and promising application prospects.
The present study delineates the developmental trajectory of defect detection technology based on machine learning techniques over time, as illustrated in Fig. 1. Notably, the delineation is anchored around the resurgence of neural networks in 2012, demarcating the overall progression of defect detection technology into two major phases: traditional machine learning algorithms and deep learning algorithms (cao2020review, ).
Figure 1. Significant technical milestones of defect detection based on machine learning.
Traditional machine learning methods include feature statistical methods, filter-based methods, and specific model methods. Defect detection methods based on feature statistics, such as Gray Level Co-occurrence Matrix (GLCM) (zhu2015yarn, ) and Local Binary Pattern (LBP) (tajeripour2007fabric, ), primarily describe defect images from a statistical perspective by analyzing the distribution of grayscale information. This approach aims to identify spatial grayscale distribution characteristics of defects for detection purposes. In addition, filter-based methods leverage signal processing principles, transforming grayscale information of defect images into the frequency domain. This allows the discovery of more easily manageable frequency domain defect features, which is particularly useful in scenarios where extracting defect features in the original grayscale domain is challenging. Specific methods include Finite Impulse Response Filter (FIR) (kumar2002defect, ), wavelet transform (li2006improving, ), Gabor filter (kumar2002defect, ), Fourier transform (chan2000fabric, ), and others. Furthermore, specific models designed for detecting surface defects in industrial products are employed, such as those based on Markov Random Field (MRF) (dogandvzic2005defect, ), Histogram of Oriented Gradients (HOG) (halfawy2014automated, ), and Deformable Parts Model (DPM) (he2019railway, ). In summary, defect detection algorithms based on traditional machine learning algorithms are well-suited for scenarios where defect features are clearly evident, defect types are limited, and there are relatively low detection requirements. These methods demonstrate a significant advantage in terms of fast detection speed.
With the resurgence of neural network technology, deep learning has experienced vibrant development and has been widely applied in defect detection, making it one of the most active research areas in the field (bhatt2021image, ). Among the various deep learning technologies relevant to defect detection, two main categories stand out. The first category involves defect detection and classification, employing techniques such as RCNN, Faster-RCNN, YOLO, SSD, etc (tulbure2022review, ; qi2020review, ). These approaches enable the detection and localization of defects in sample images, providing information on the defect’s category. The second category involves defect segmentation, utilizing algorithms like FCN, U-Net, PSPNet, DeepLab, etc (tabernik2020segmentation, ). These segmentation algorithms not only detect defects but also offer pixel-level classification and segmentation. However, compared to defect detection and classification algorithms, pixel-level segmentation algorithms are more time-consuming. It is noteworthy that defect detection solutions based on deep learning technologies can meet the majority of modern defect detection requirements. Unlike traditional machine learning algorithms that require manually designed defect features, deep learning algorithms can automatically extract defect features. Moreover, these detection algorithms demonstrate robustness in handling multiple defect types, complex backgrounds, and large-scale detection tasks. However, a drawback is that training detection models demand substantial resources, including computational power and the collection and annotation of massive training data.
2.3. Material Defect Detection
Material defect detection, as an important branch of defect detection, is a crucial area of research in industrial production to prevent safety issues caused by material defects. Rapid and non-destructive detection of potential defects in the material production process is a critical aspect of ensuring material quality. Common material defects can be categorized into surface defects and structural defects. Surface defects typically manifest as localized irregularities on the material surface, including irregular or abnormal features such as cracks, scratches, dents, colour variations, bubbles, wear, foreign objects, and more (zhao2020hole, ; raizada2021surface, ). Structural defects, on the other hand, involve abnormalities in the overall structure and shape of the material, such as shape defects (deformation, twisting, or bending), structural defects (improper connections, joints, or fixations), structural inconsistencies, and defects in connecting parts. Structural defects often exhibit a more complex manifestation in real-world applications and may include surface defects (zhao2023structural, ; li2020strong, ; zhang2020defect, ). Fig. 5, Fig. 5, Fig. 5, and Fig. 5 illustrate some common defect samples found in materials, where the Image column contains the original defect images and the GT column includes the corresponding ground truth labels.
Common methods for material defect detection generally include magnetic particle testing (MPT), penetrant testing (PT), eddy current testing (ECT), ultrasonic testing, X-ray inspection, and machine vision inspection. The relevant introduction and comparison among these detection methods are as below:
  • MPT is a non-destructive testing method used to detect surface and near-surface defects in materials. It relies on the properties of magnetic powders, which are applied to the material surface after the application of a magnetic field. If defects are present, the magnetic particles gather at the defect locations, forming visible lines or spots, thereby indicating the position and shape of the defect. MPT is sensitive and cost-effective, but it is limited to detecting surface defects in ferromagnetic materials and may not be suitable for real-time automated detection (burke2013review, ; lugin2023automated, ).
  • PT is a non-destructive testing method for discovering surface or near-surface cracks, pores, or other open defects. It relies on the penetration behaviour of a liquid into the defect surfaces, revealing hidden flaws in the tested object. PT is applicable to various materials, offering advantages in surface defect detection, but it may have limitations in detecting defects in porous materials, and the detection speed is relatively slow with sensitivity influenced by various factors (kumar2015review, ; mehta2015surface, ).
  • ECT is a non-destructive testing method for detecting surface and near-surface defects in conductive materials. It is based on Faraday’s electromagnetic induction principle, generating eddy currents by introducing an alternating magnetic field to identify defects through variations in eddy currents. ECT is fast, sensitive, and suitable for high-temperature defect detection. However, it may lack accuracy in displaying defect shapes and sizes, and the detection progress can be slow (hughes2014near, ).
  • X-ray Inspection is a non-destructive testing method utilizing X-rays to penetrate materials and measure the material’s absorption of X-rays, providing information about internal structures and defects. This method is versatile, and capable of detecting both surface and internal defects with high efficiency. However, it poses radiation risks to operators, and the equipment is generally complex and expensive, with detection accuracy influenced by environmental conditions (du2019approaches, ).
  • Ultrasonic Testing is a non-destructive testing method using ultrasonic waves to detect internal or surface defects in materials. It is based on the propagation characteristics of ultrasonic waves, sending pulses and receiving reflected signals to identify abnormalities in the material. Ultrasonic testing is simple, efficient, and applicable to both surface and internal defect detection. However, it can be more challenging for complex architectural materials, and its effectiveness is limited for non-conductive materials (honarvar2020review, ).
  • Machine Vision Inspection is a technique utilizing computer vision systems for defect detection and quality control. It involves capturing images and using computer algorithms to analyze and process images to detect, recognize, and evaluate defects in the tested object. Machine vision is versatile, highly accurate, and not significantly influenced by material and defect types, making it suitable for various applications (ren2022state, ).
Figure 2. Textile fabric defects in (silvestre2019public, )
Figure 3. Magnetic tile surface defects in (huang2020surface, )
Figure 4. Commutator surface defects in (Tabernik2019JIM, )
Figure 5. Textured surface defects in (wieler2007weakly, )
Through the above comparison of existing methods for material defect detection, it is evident that machine vision-based defect detection methods have significant advantages compared to other approaches. They enable automated non-destructive testing for virtually any material defect, offering high efficiency, safety, and precision. Importantly, defect detection through machine vision seamlessly integrates with the rapidly evolving field of artificial intelligence, making existing material defect inspection more intelligent. This approach has garnered significant attention and preference from both academia and industry. Therefore, machine vision-based defect detection methods will serve as the central focus of this paper.
The machine vision defect detection system consists of four main components: image acquisition, image preprocessing, image analysis, and defect information output and recording, as illustrated in Fig. 6. In the image acquisition phase, a visual system composed of lenses, lighting sources, industrial cameras, etc., is established to capture visual information of material defects in the form of image streams. After image acquisition, preliminary image preprocessing is required on the raw image information, including conventional operations such as noise reduction, image enhancement, and region of interest extraction, to prepare for subsequent image defect analysis.
The image analysis phase constitutes the core of the entire machine vision defect detection system, as specific defect detection algorithms will be deployed and executed here. For defect detection using traditional unsupervised machine learning algorithms, manually designed defect feature extraction and classification algorithms can be directly executed. If deep learning algorithms are employed for defect detection, steps such as defect image annotation, neural network model design, and model training are necessary, ultimately resulting in a defect detection model that can be deployed and executed. Lastly, the visual detection system records defect feature information outputted by the defect detection model, such as defect category, defect size, defect location, etc., in a database. Necessary actions are then taken for the defective material, such as alarms or removal.
Figure 6. Illustration of material defect detection system based on machine vision.
2.4. Challenges in Material Defect Detection
Compared to common defect detection tasks, material defect detection technology necessitates consideration of specific industry attributes, material characteristics, and the material production environment. Consequently, this introduces new challenges and complexities to the realm of material defect detection. The current challenges facing machine learning-based material defect detection include:
  • Sample Scarcity: Sample scarcity is a common challenge faced by all current defect detection systems, primarily manifested in three aspects: (1) With the advancement of modern production processes, strict control over the yield of substandard products has become a norm in current material production. This results in a scarcity of defect samples compared to readily available normal samples. (2) The collection and annotation of defect sample data incur high costs, rendering the acquisition of high-quality defect samples even more challenging. (3) Due to factors such as business competition and privacy protection, defect samples are rarely disclosed publicly, making it difficult to access actual defect samples.
  • Data Quality Issues: Data quality is crucial for the performance of machine learning models. Low-quality data may introduce noise, leading to misjudgments or missed detections. Improving data quality is a significant challenge in material defect detection using machine learning, as the task requires a substantial amount of annotated data for training. However, practical applications often encounter issues such as data fuzziness and noise interference, impacting model training effectiveness and detection performance.
  • Identification of Complex Defects: Recognizing complex defects poses a significant challenge in machine learning for material defect detection. Efficient feature extraction and classification algorithms are required to accurately identify these defects, extracting useful features to enhance detection performance. The diverse types, shapes, sizes, and positions of defects in materials make identification challenging. Traditional detection methods often struggle to accurately identify complex defects such as cracks and pores.
  • Multimodal Data Processing: Handling multimodal data is a crucial challenge in machine learning for material defect detection. Different data types exhibit distinct characteristics and noise levels, necessitating the development of specialized techniques to address these differences. Additionally, integrating data from various sources to improve detection performance presents a challenge. Material defect detection involves multiple data types, including images, ultrasonic waves, X-rays, etc. Effectively processing these multimodal data, extracting useful features, and enhancing detection performance pose challenges.
  • Real-Time Online Detection: Real-time online detection is a crucial challenge in machine learning for material defect detection. Developing efficient algorithms and optimizing models to reduce computation time and resource consumption are necessary to achieve real-time online detection. Maintaining model performance and accuracy in real-time environments is also challenging. Fast, reliable, and real-time defect detection is essential on production lines or in real-time applications, but some current machine learning methods may struggle to meet these requirements.
  • Interpretability and Transparency: Interpretability and transparency are important challenges in machine learning for material defect detection. Developing new technologies and methods to analyze the internal workings of models is essential to enhance model transparency and interpretability. Striking a balance between high performance and improved model transparency and interpretability is a challenge. For critical application areas like medical devices and safety-related fields, increasing model transparency and interpretability is a crucial requirement. However, current deep learning models are often perceived as ”black boxes,” making it difficult to explain the reasons behind their decisions.
  • Privacy and Security: Privacy and security present important challenges in machine learning for material defect detection. Efficient data encryption techniques and secure storage solutions need to be developed to protect privacy and security. Additionally, ensuring models are not subject to malicious attacks or tampering is essential. Privacy and security issues must be considered when dealing with material defect detection. Challenges include secure data storage and transmission, as well as maintaining the confidentiality of models.
2.5. Survey Overview
Facing numerous challenges in material defect detection, the rapid development of visual hardware and computer vision technologies, particularly those based on deep learning, has led machine learning-guided defect detection algorithms to gradually dominate research. Researchers have innovatively applied various machine learning algorithms to defect detection, significantly improving accuracy, efficiency, and applicability. This study, centred around defect detection technology, explores the application scenarios and advantages and disadvantages of different material defect technologies. It not only provides a comprehensive literature review and future research directions for subsequent researchers but also offers extensive guidance for industrial practitioners.
This paper summarizes recent hot research technologies applied to material defect detection in Fig. 7 and provides detailed discussion and analysis in the following content. Unsupervised detection algorithms are further categorized into traditional unsupervised machine learning algorithms and unsupervised deep learning algorithms. For situations where defects are relatively simple and defect types are known, handcrafted defect features can be used, and rapid defect detection can be achieved using traditional unsupervised machine learning algorithms. For defects without manual annotations, unsupervised deep learning algorithms are needed to automatically learn defect features from defect samples and accurately differentiate between normal and defective samples. Additionally, in cases where a sufficient number of annotated defect samples are available, supervised learning algorithms can be employed. Generally, deep neural network models are used for self-learning features, enabling defect detection, classification, and segmentation. However, in realistic industrial scenarios, annotated defect samples are often scarce. Thus, the latest semi-supervised or weakly supervised learning algorithms are applied to make reasonable learning decisions for labelled and unlabeled defect samples. In situations where defect samples are extremely scarce or the ratio of defect to normal samples is highly imbalanced, modern generative learning methods can be applied for the artificial augmentation of defect samples, balancing the disparity in the quantity of positive and negative samples. For learning new defect samples, reinforcement techniques can also be applied. Furthermore, to meet the industrial detection criteria for real-time and high detection efficiency, model compression techniques are actively applied during the deployment stage to enhance the operational efficiency of defect detection models. Although machine learning-based material defect detection has achieved overall improvements in detection accuracy and efficiency, challenges such as the complexity of defect types, generalization issues in defect detection models, and high costs in model training and deployment still exist.
Figure 7. Common machine learning methods for conducting the material defect detection.
3. Techniques for Material Defect Detection
3.1. Unsupervised Learning Techniques in Material Defect Detection
This section elaborates on the unsupervised learning techniques employed in material defect detection. Fig. 8 outlines its general workflow comprising four main steps: data collection, data preprocessing, feature extraction, and anomaly detection. Initially, a sensor captures data, forming the original dataset. This dataset is then preprocessed by removing outliers, normalizing values, and filling in missing data to produce a processed dataset. Key features are extracted using methods such as PCA, histograms, and DBNs. Finally, anomaly detection is performed using techniques like clustering, SVM, and classification networks to identify and classify material defects.
Figure 8. Illustration of the general working process for conducting the material defect detection with unsupervised learning techniques.
3.1.1. Histogram-based Techniques
Histogram-based techniques have widely been used in material defect detection (kumar2021histogram, ). The method involves dividing a wood surface image into equal-sized blocks, analyzing each block’s colour using a percentile colour histogram, and extracting texture features through singular value decomposition (SVD) (song2015wood, ). These features are combined into a feature vector for each block. A support vector machine (SVM) classifier is then used to identify defects, particularly effectively detecting knot-type defects on wood surfaces. An improved Otsu method, Weighted Object Variance (WOV) (yuan2015improved, ), is proposed for efficient surface defect detection in machine vision. WOV adjusts the thresholding technique, traditionally used in image processing, to better handle varying defect sizes and histogram types. It ensures more accurate segmentation, leading to higher defect detection rates and lower false alarms. Experimental results demonstrate WOV’s superiority over traditional methods like Otsu and maximum entropy in detecting defects on various surfaces. Erazo-Aux et al. (erazo2019histograms, ) introduce a new method for detecting defects in thermal images of composite materials using pulsed thermography. It uses local histograms of oriented gradients for simple yet effective differentiation of defects from the material background without relying on image enhancement or complex analysis. The method demonstrates high accuracy, particularly in images of carbon fibre-reinforced plastic, achieving robust detection with an average AUC score of 0.949. This approach is more straightforward and less costly than other methods requiring extensive preprocessing. In order to detect the defects in sewers, Mahmoud et al. (halfawy2014automated, ) present a pattern recognition algorithm designed to automate defect detection and classification in sewer pipes using CCTV inspection images. It employs histograms of oriented gradients (HOG) and a support vector machine (SVM). The process involves image segmentation to identify potential defect areas, followed by SVM classification using trained HOG features.
3.1.2. Clustering Algorithm
Clustering algorithms are a class of unsupervised machine-learning techniques that can effectively detect material defects (jin2020machine, ). These algorithms group similar data points together based on certain features, and this characteristic makes them suitable for identifying anomalies or defects in materials. Thomas et al. (leung2001representing, ) focus on and develop a method to identify materials like concrete, rug, marble, or leather based on their textures, focusing on surface reflectance and normals. We introduce ”3D textons,” small prototype surface patches with unique geometric and photometric properties, such as ridges or grooves. Each texton is associated with an appearance vector, capturing light distribution under various conditions. Using a clustering technique, we create a compact library of these textons. This allows for the characterization and recognition of materials from a few images, even under novel lighting and viewing conditions. The method also predicts how materials appear in different environments. In high-tech materials such as semiconductor manufacturing, a spatial defect diagnosis system for semiconductor manufacturing has been proposed to estimate defect clusters and their separation, including both convex and non-convex types. The approach (wang2008recognition, ) uses a spatial filter to clean wafer maps and a hybrid of entropy fuzzy c-means and spectral clustering to isolate defect patterns. A decision tree based on cluster features helps identify defect types. Tested with data from a DRAM company in Taiwan, the method effectively extracts and classifies various defect patterns, showing potential for application in other industries like liquid crystal and plasma display manufacturing. For steel materials, unsupervised machine learning was used to classify images from the Northeastern University Steel Surface Defects Database, which contains micrographs of different defects in hot-rolled steel (cohn2021unsupervised, ). The study employs the VGG16 convolutional neural network, pre-trained on the ImageNet dataset, to extract features from the micrographs. The principal component analysis is then applied to refine these features, followed by k-means clustering for image classification without requiring labelled data. The method achieves an impressive 99.4% accuracy and can classify new images without retraining. This represents a significant improvement over previous studies. Additionally, a sensitivity analysis is conducted to understand the impact of each step on the classification performance, offering valuable insights for applying unsupervised machine learning in materials science.
Unlike K-means clustering, Hierarchical Clustering builds a hierarchy of clusters through a bottom-up approach (agglomerative) or top-down approach (divisive). Jorge et al. (igual2020hierarchical, ) presents a hierarchical clustering approach using impact-echo techniques for defect detection in materials. Feature vectors, indicating wave propagation changes caused by defects, are analyzed as mixtures of Gaussians. Clusters are merged using the Kullback-Leibler divergence. Applied to aluminium alloy specimens, the method successfully classifies them into categories like no defect, one hole, one crack, and multiple defects, effectively distinguishing between defective and non-defective materials. In addition, Brzakovic et al. (brzakovic1990approach, ) introduce an expert system for identifying and categorizing defects in materials with complex textures using digitized images. It involves a two-step image segmentation process and a material-specific defect classification approach. A hierarchical classification scheme incorporating shape, size, and texture descriptors is used for wood. The system, tested on parquet samples, has successfully detected and classified various defects.
3.1.3. Principal Component Analysis
Principal Component Analysis (PCA) is a statistical technique used in material defect detection to simplify the complexity of high-dimensional data while retaining trends and patterns (cao2024crack, ). Active thermography, a nondestructive testing method for composite materials, is improved in this study with a Sparse Principal Component Thermography (SPCT) method (wu2018sparse, ). Building on the advantages of Principal Component Thermography (PCT), SPCT introduces a penalization term, enhancing flexibility and interpretability due to its structural sparsity. The effectiveness of SPCT is proven through experiments on a carbon fibre-reinforced plastic specimen, demonstrating superior subsurface defect visualization compared to PCT. Machine learning methods are increasingly significant in the field of nondestructive testing, particularly for the quality assessment of polymer composites. Among these, a Generative Adversarial Network (GAN), a form of deep learning, is introduced to thermography as an image augmentation tool, enhancing defect detection capabilities (liu2020generative, ). This method uses data augmentation to produce more diverse and informative images, improving the defect detection capabilities of thermographic analysis. The effectiveness of GPCT is demonstrated through successful experiments on a carbon fibre-reinforced polymer specimen, showcasing its potential in nondestructive testing.
3.1.4. Self-organizing Maps
Self-organizing Maps (SOMs) are a type of unsupervised neural network used for data visualization and pattern recognition, which can be effectively applied to material defect detection (shakhovska2020improved, ). An automated visual inspection method for detecting and classifying textile defects, using the autocorrelation function for feature extraction and a Self-Organizing Feature Map (SOFM) (tolba1997self, ). This approach is immune to illumination changes and noise, thanks to the noise rejection property of the autocorrelation function. Tested on real textile samples, the SOFM effectively clusters the feature vectors, showcasing the advantages of unsupervised learning systems in defect classification. For textured pavement images, Mathavan et al. (mathavan2015use, ) study a method using a self-organizing map (SOM) for detecting road cracks in textured pavement images. It involves analyzing images segmented into cells’ texture and colour properties and using a Kohonen map to identify cracks. The technique achieves 77% precision and 73% recall, showing effectiveness even in uncontrolled environments and outperforming an advanced algorithm in reducing false positives in crack detection.
3.1.5. Variational Autoencoders
Variational Autoencoders (VAEs) are a type of deep learning model particularly well-suited for anomaly and defect detection in materials. Using unsupervised learning, Tsai et al. (tsai2021autoencoder, ) evaluate an improved Convolutional Autoencoder (CAE) model for automated defect detection in manufacturing. Unlike traditional CAEs that minimize reconstruction errors, this enhanced CAE incorporates a regularization to tighten the feature distribution of defect-free samples. This approach makes defect samples stand out more distinctly during evaluation. Tested on various material surfaces, including textured and patterned ones, the modified CAE demonstrates superior performance in defect detection compared to conventional CAEs. In addition, a variational autoencoder-enhanced deep learning model (VAEDLM) for efficiently classifying imbalanced wafer defects in semiconductor foundries (wang2021variational, ). The model effectively generates and recognizes wafer defect patterns by utilizing variational autoencoders and a refined convolutional neural network. Tested on the WM-811K dataset, the VAEDLM achieves exceptional performance metrics including 99.19% accuracy and 99.96% AUC, outperforming current methods. Additionally, the model’s interpretability is enhanced through saliency maps and t-SNE, addressing a common limitation in deep learning applications.
3.1.6. Deep Belief Network
A Deep Belief Network (DBN) is a type of deep neural network with multiple layers of stochastic, latent variables. It’s particularly useful for feature extraction and classification tasks, making it suitable for material defect detection. Cang et al. (cang2017microstructure, ) focus on enhancing Integrated Computational Materials Engineering (ICME) by introducing a Convolutional Deep Belief Network (CDBN) for feature learning. This approach automates the transformation of microstructures into lower-dimensional representations, achieving a significant reduction in dimensionality. The model is applied to diverse material systems, including alloys and sandstone, successfully reconstructing materials closely resembling the originals in terms of key properties. This method outperforms traditional techniques that rely on the Markovian assumption, showcasing its potential in advanced material design and reconstruction. Furthermore, Liu et al. (liu2019real, ) introduce a real-time quality monitoring and diagnosis scheme for manufacturing processes using a DBN. It addresses the challenge of handling complex, high-dimensional process data by converting them into quality spectra, which are then analyzed using a DBN. This model is developed through offline learning and applied for online process monitoring and diagnosis. Its effectiveness is demonstrated through simulation experiments and a practical example in injection moulding, showing superior performance compared to traditional methods.
3.1.7. Unsupervised Technique Summary
The application of unsupervised learning techniques in material defect detection exhibits diversity and versatility. Table 2 summarizes the application scenarios, advantages, and limitations of various unsupervised techniques. As shown in Table 2, Histogram-based techniques are applicable for detecting specific types of defects, such as nodular defects on wood surfaces, but may not be suitable for other types of defects. Clustering algorithms are capable of detecting material defects by grouping similar data points, offering wide applicability for identifying anomalies or defects, albeit potentially requiring a large amount of unlabeled data to enhance accuracy. PCA simplifies the complexity of high-dimensional data, yet may be insufficient for handling more complex defect types. SOMs are utilized for data visualization and pattern recognition, effectively classifying feature vectors, but may yield suboptimal results outside controlled environments. VAEs highlight defect samples and are suitable for detecting anomalies and defects in materials, yet may necessitate a significant amount of training data. DBNs automatically transform microscopic structures and are employed for feature extraction and classification tasks, though they may require substantial training and computational resources. Considering task requirements, data characteristics, and computational resources, selecting the appropriate technique is crucial for improving the accuracy and efficiency of material defect detection.
Table 2. Unsupervised Technique Comparison in Material Defect Detection
Unsupervised Technique Application Scenarios Advantages Limitations
Histogram-based techniques
specific types of defects, such as
nodular defects on wood surfaces
effective identification of specific
types of defects
not applicable to
other types of defects
Clustering algorithms multiple similar detects identify generic anomalies or defects
require a large amount of
unlabeled data
Principal Component Analysis
simplify the complexity of
high-dimensional
detect feature data
retain defect trends and patterns
not sufficient to handle more
complex defect types
Self-organizing Maps
defect data visualization
and pattern recognition
effectively classify detect
feature vectors and
insensitive to the distribution
and features of defect data
highly sensitive to parameter
selection and high
computational complexity
Variational Autoencoders
Scarcity of defect samples,
anomaly recognition
in defect detection.
enrich training samples of defects
and identify anomalies
in defect detection.
require a large amount
of training data
Deep Belief Network
defect feature extraction
and classification
identify anomalies
in defect detection.
require significant training
and computational resources
3.2. Supervised Learning Techniques in Material Defect Detection
This section summarizes the supervised learning techniques relevant to material defect detection. Fig. 9 outlines a general workflow for this process, divided into five main steps: data collection, data labelling, data preprocessing, model selection and training, and model testing. Initially, a sensor captures data to form the original dataset. This dataset is then labelled with defect categories in the data labelling step. Next, the data undergoes preprocessing, which includes data cleaning, augmentation, and normalization, resulting in a processed dataset. In the model selection and training step, a designed learning model is trained using forward and backward propagation methods on the processed dataset. Finally, the well-trained model is tested with test samples to detect and classify material defects.
Figure 9. Illustration of the general working process for conducting the material defect detection with supervised learning techniques.
3.2.1. Regression Algorithms
Regression algorithms are a class of statistical methods used in machine learning for predicting a continuous outcome. They can be effectively applied to material defect detection, particularly in predicting the extent, severity, or type of defects based on various input features. To detect the defection efficiently, a two-stage neural algorithm for defect detection and depth characterization using active thermography is proposed (dudzik2015two, ). The first stage involves a classification neural network for detecting defects, followed by a regression neural network to estimate their depth. Results from experimental investigations and simulations are presented, along with a sensitivity analysis that examines the effects of emissivity and ambient temperature errors on depth estimation. The algorithm’s performance is validated using a test sample of low thermal diffusivity material. Various regression algorithms have been utilized in detecting defects during the Laser Beam Additive Manufacturing (LBAM) process. In their work, Mahmoudi et al. developed a system for detecting anomalies specifically in the Laser Powder Bed Fusion (LPBF) printing process (mahmoudi2019layerwise, ).
3.2.2. Neural Network Algorithms
For a long time, the practical use of Thermographic Non-Destructive Evaluation (TNDE) has faced challenges due to the intricate and nonlinear nature of related inversion problems, such as accurately determining the depth of defects. However, artificial neural networks have recently shown promising results in detecting and quantitatively assessing defects in TNDE. Darabi et al. (darabi2002neural, ) introduce a three-dimensional thermal model for non-homogeneous materials, like carbon fibre-reinforced plastic (CFRP). The outcomes of this model are then compared with analytical solutions derived from Duhamel’s theorem. Subsequently, the paper describes the development of two back-propagation neural networks, one for detecting defects and the other for estimating their depth. Lastly, it presents and examines both simulated and experimental results. Furthermore, to improve the detection performance, an advanced convolutional neural network-based method for accurately detecting surface defects in workpieces has been proposed (xing2021convolutional, ). It introduces a specialized model (SCN) for feature extraction and uses an optimized IOU (XIoU) for loss function definition. The method demonstrates superior performance over other models in both classification and detection tasks, with high accuracy and real-time detection capabilities, making it highly effective for automatic surface defect detection in workpieces.
3.2.3. Ensemble Learning Algorithms
Ensemble learning in material defect detection involves using multiple learning algorithms or models to obtain better predictive performance than could be obtained from any of the individual models alone (alipour2020increasing, ). A novel method for robotic hammering inspection aimed at maintaining social infrastructures, with a specific focus on estimating material defects like delamination depth in concrete (fujii2016defect, ). With the aging of infrastructure, there’s a growing need for automated diagnostic methods. The hammering test, known for its efficiency in inspections, is gaining interest in robotic automation. The study introduces a method to estimate defect conditions suitable for detailed robotic inspections. It also incorporates an integration technique of multiple classifiers to enhance inspection accuracy. Additionally, the system is designed to minimize the impact of mechanical running noise. Experimental results with concrete test samples show the effectiveness of this approach, confirming its accuracy in both defect detection and condition estimation. In addition, a new deep ensemble learning method for real-time surface defect inspection in automated production focuses on overcoming challenges like domain shifts and imbalanced training data in flowing data scenarios (liu2019data, ). It introduces a unique distribution discrepancy identifier that blends deep CNNs with shallow learning for improved robustness and precision. The method was successfully tested on an industrial surface mount technology line, showing excellent mean average precision and adaptability for streaming data in industrial settings.
3.2.4. Attention Mechanisms Related Algorithms
Attention mechanisms in material defect detection refer to a set of techniques in machine learning, particularly in deep learning, that enable the model to focus on specific parts of the input data that are more relevant to the task at hand. Fang et al. (fang2022tactile, ) propose a novel fabric defect detection method using a vision-based tactile sensor, which is more robust than traditional visual methods as it’s unaffected by dyeing patterns and ambient light. A robotic arm with tactile sensors automates data collection, creating comprehensive fabric datasets. The approach employs a convolutional neural network (CNN) with an attention mechanism and frequency domain filtering to enhance defect detection accuracy and efficiency. Experiments demonstrate its superiority over visual defect detection methods, confirming its practicality and efficiency for real-world applications. In addition, Chen et al. (chen2021defect, ) present a method for detecting defects in aluminium profile surfaces (APS) using a deep self-attention mechanism (DSAM) specifically designed to work effectively in noisy image conditions. This method improves upon traditional approaches by incorporating a residual learning strategy for better feature extraction from noisy images, employing a self-attention mechanism to refine these features, and using a softmax classification layer to identify defects in APS accurately. The method has achieved higher accuracy in detecting APS defects under various noise conditions.
3.2.5. Supervised Technique Summary
This section summarizes the application of supervised learning techniques in material defect detection, including regression algorithms, neural network algorithms, ensemble learning algorithms, and attention mechanism-related algorithms, with a comparison provided in Table 3. Regression algorithms are suitable for predicting the severity and extent of defects, offering high detection accuracy, but requiring a large amount of labelled data and difficult adaptation to new types of defects. Neural network algorithms are applicable to complex and nonlinear problems, enabling accurate detection and quantitative assessment of defects, but demanding substantial training data and computational resources. Ensemble learning algorithms improve predictive performance by combining multiple models, enhancing detection accuracy and robustness, although the models may become overly complex and challenging to interpret. Attention mechanism-related algorithms enable models to focus on task-relevant parts, enhancing the accuracy and efficiency of defect detection, yet requiring adjustment and optimization to suit specific application scenarios. Considering task requirements, data characteristics, and computational resources, selecting the appropriate techniques is crucial for improving the accuracy and efficiency of material defect detection.
Table 3. Supervised Technique Comparison in Material Defect Detection
Supervised Technique Application Scenarios Advantages Limitations
Regression
Various types of material defects,
such as surface defects,
internal defects, etc.
accurate predictions of defectiveness
and severity, as well as quantitative
assessments of defects.
need Large amounts of labeled data
and be hard to adapt to new defects
Neural Network
Complex defect feature extraction and
classification with intricate
backgrounds, structures, and noise.
Strong adaptability to different
defect types, capable of learning
complex defect features and
accommodating large-scale
defect detection.
Requires extensive annotated defect
data, prone to model overfitting and
black-box decision-making.
Ensemble Learning
Multimodal defect types,
complex defect detection task
Enhanced defect detection performance
and model generalization, high flexibility
High algorithm complexity, high
hardware infrastructure requirements,
demanding defect data quality
Attention mechanisms
Detection of defects of
various types and scales
Enhanced defect detection accuracy
and efficiency by focusing on local
defect features, simultaneously
improving model adaptability
Increased computational complexity,
increased defect data volume and
annotation costs, poor model interpretability
3.3. Semi-supervised Learning Algorithms
In practical industrial scenarios, the collection and annotation costs of defective samples are often high, leading to a significant amount of unlabeled data. In order to make more effective use of these unlabeled data, researchers have turned to the setting of semi-supervised learning (liu2022defect, ; liu2021semisupervised, ; wang2021new, ). In semi-supervised learning, the training set includes both normal and abnormal samples, but only a small portion of the samples are labelled. Compared to supervised learning, semi-supervised learning reduces the dependence on labelled data.
[…]
Compared to unsupervised methods, semi-supervised learning significantly improves model performance when provided with a small amount of labelled defective samples. This approach is beneficial in addressing new types of defects in industrial product manufacturing. The application of semi-supervised learning is primarily focused on defect classification at the image level, while research on defect localization tasks is still an area that requires further exploration.
3.4. Reinforcement Learning Algorithms
Material defect detection often faces complex and dynamic environments, where the types and features of defects may exhibit diversity. Reinforcement learning can learn optimal strategies through interaction with the environment, enhancing the effectiveness of material defect detection. Therefore, reinforcement learning, capable of meeting the demands of more complex material defect detection environments and achieving autonomous real-time dynamic learning of material defects, has started to gain favour among researchers.
In (chu2020neural, ), a reinforcement learning-based neural batch sampler is employed to select training batches for an autoencoder that are more conducive to generating loss curves distinguishing abnormal and normal regions. Additionally, periodic reinitialization and retraining of the autoencoder allow it to start from different initial weights and optimize towards different local minima. Throughout the training process, reinforcement learning gradually reduces its dependence on behaviour cloning rewards while increasing reliance on prediction loss rewards, facilitating a smooth transition from a heuristic policy to a more optimized strategy. In addition, (chung2022reinforcement, ) introduces a reinforcement learning-based online defect detection and mitigation method to enhance quality assurance in additive manufacturing processes. The proposed algorithm, called Continuous G-Learning, simultaneously utilizes two types of prior knowledge (offline and online) to reduce the number of learning samples by adjusting process parameters, ensuring the learned policy is optimal. Experimental results demonstrate the ability of this method to rapidly detect and mitigate emerging defects, improving surface quality and mechanical performance of printed products. Also, In (chung2022reinforcement, ), a reinforcement learning-based virtual plastic texture surface defect generation system is proposed for data augmentation in deep learning detection systems. The reinforcement learning network, consisting of an actor network and a critic network, searches for optimal rendering and defect parameters to generate realistic texture images. Using the generated images to supplement the original training dataset enhances data diversity and balance, improving the detection system’s recall rate. Lastly, (zhang2023duak, ) collects and integrates heterogeneous steel surface defect data from a human-network-physical system, constructing a steel surface defect knowledge graph. The proposed bidirectional reinforcement learning-based knowledge graph inference method employs two agents exploring paths from opposite directions, leveraging the historical experience and path-sharing mechanisms to update action selection policies and reward functions, enhancing inference efficiency and accuracy. Applied to steel surface defect detection and diagnosis, this method leverages knowledge graph inference to supplement missing knowledge, providing decision support for process improvement. In (mohanty2024enhancing, ), a digital twin model of a titanium alloy washer is created using Blender and Blendtorch, incorporating geometry, surface texture, lighting conditions, and camera parameters. A Proximal Policy Optimization (PPO) algorithm is employed for reinforcement learning to train an agent, interacting with the Blender environment to learn how to adjust rendering parameters to deceive the discriminator and enhance authenticity. A defect detector based on ResNet18 is trained using synthetic defective and non-defective washer images and evaluated for performance and robustness on real washer images.
Reinforcement learning methods can adapt to continuously changing environments in material defect detection. This is because they learn through interaction with the environment, enabling them to continuously adjust their strategies during training to accommodate new defect types or environmental changes. They also exhibit high real-time capability, allowing learning and decision-making in real-time data streams, which is crucial for scenarios requiring timely responses in material defect detection. Additionally, reinforcement learning models can autonomously learn defect features without the need for extensive manual feature engineering. This enables them to extract valuable information from data and exhibit adaptability to complex defect situations. However, reinforcement learning typically demands substantial training and computational resources, especially when dealing with complex material defect detection tasks. This may result in long training times and high computational costs. Moreover, reinforcement learning models usually require a large number of samples for training, particularly in the case of deep reinforcement learning models. Acquiring large-scale annotated data in the field of material defect detection can be an expensive and challenging task. Overall, reinforcement learning holds potential in material defect detection.
3.5. Generative Learning Algorithms
In the industrial domain, the lack of precisely annotated defect samples makes it challenging to effectively train neural network models. To overcome this issue, data augmentation becomes a crucial factor in enhancing model performance. Common image augmentation techniques include cropping, rotation, distortion, and affine transformations (zhou2023comparative, ; huang2020surface, ; dabetwar2022fatigue, ). However, these methods, while altering the shape of defects, still maintain defects within the original background and may lack sufficient diversity, potentially leading to image distortion problems.
To obtain more diverse augmentation data, generative learning techniques are introduced into the augmentation of defect sample data, aiming to synthetically create realistic defects on normal samples. The work by (valente2020print, ) introduces an end-to-end deep learning framework based on semantic segmentation for mapping printing defects. This framework enables pixel-level detection and localization of defects, such as stripes and colour bands, in printed images. Two types of printing defects and printing scan effects are simulated using image processing and computer graphics techniques, generating synthetic training datasets. This approach avoids the costs and time associated with manually collecting and annotating data. Experimental results demonstrate that the proposed framework can effectively learn features from synthetic data and generalize them to real printing defects. In (bosse2023automated, ), a data-driven feature labelling model is proposed for detecting hidden defects, such as pores in aluminium alloy die-cast plates, in low-cost X-ray projection images. Due to the absence of accurate annotated data for real X-ray images, the authors use numerical simulations of X-ray images to generate synthetic training datasets, taking into account various noise and distortion factors. Furthermore, (farady2023preaugnet, ) introduces an offline PreAugmentation Network (PreAugNet) for generating and filtering augmented samples to improve the training of target networks. The PreAugmentation Network utilizes a convolutional neural network (CNN) to extract features from original samples and updates the generated samples through the decision boundary analysis of a support vector machine (SVM) classifier. This ensures that the new samples produced by the PreAugmentation Network are more diverse and representative. Evaluation on three real industrial steel surface defect datasets demonstrates the effectiveness of PreAugmentation Network data augmentation in improving the training of defect classification networks.
In addition to the methods mentioned above for generating defect data, the most common generative techniques are autoencoders and Generative Adversarial Networks (GANs). (yun2020automated, ) proposes a metal surface defect classification algorithm based on Deep Convolutional Neural Networks (DCNN), utilizing metal surface images collected by a machine vision system as input. To address the issue of imbalanced defect data, they introduce a data augmentation approach based on the Conditional Convolutional Variational Autoencoder (CCVAE). By learning the distribution of defect data, CCVAE generates images of various defect classes. Training the DCNN classification model on data generated by CCVAE enhances the model’s generalization performance and accuracy. Besides autoencoders, GANs is the most popular technique employed to generate the defect images. (pourreza2021g2d, ) presents a defect detection method based on Generative Adversarial Networks (GAN), named G2D (Generate to Detect Anomaly), for detecting anomalies in images and videos with only normal samples. G2D leverages the generator of GAN to produce different types of anomaly samples during training, including randomly generated samples deviating from the normal distribution and semantically inconsistent samples close to the boundary. (zhang2021defect, ) introduces Defect-GAN, a defect sample synthesis method based on Generative Adversarial Networks (GAN), to address the issue of insufficient defect samples in automatic defect detection. Defect-GAN simulates the damage and restoration process on a large number of normal samples to generate realistic and diverse defect samples. (jain2022synthetic, ) proposes a data augmentation method based on Generative Adversarial Networks (GAN) to improve the performance of surface defect detection and classification. Three different GAN architectures (DCGAN, AC-GAN, and InfoGAN) are employed to generate surface defect images containing six types of hot-rolled steel strip defects. Additionally, the study explores the latent space of GANs, discovering that adjusting continuous or discrete latent variables can control certain features of synthetic images, such as the severity of defects and lighting conditions, thereby incorporating additional domain knowledge.
In summary, generative learning methods can be employed in material defect detection to synthesize defect samples, thereby increasing the diversity of training data. This is particularly beneficial for material defect detection tasks that lack a large number of annotated defect samples, as it can significantly alleviate the cost of defect annotation, address data imbalance issues, and enhance model robustness. However, the performance of generative learning methods largely depends on the quality of the generated model, which may be constrained by the limitations of training data and may struggle to fully replicate the real-world distribution of defects. Additionally, synthesized defect samples may not be sufficiently diverse, leading to a decline in model performance in practical applications. Moreover, some generative learning methods, especially those based on deep learning, may have high computational resource requirements, especially when training and generating on a large scale. Therefore, the choice of generative learning methods should be comprehensively considered based on specific task requirements, data characteristics, and computational resources.
3.6. Comparison for Diverse Material Defect Detection Tasks
Table 4. Comparison for diverse material defect detection tasks.
Task Setting Applicable Scenarios Advantages Limitations
Unsupervised
Learning
$\bm{*}$ No labeled training data.
$\bm{*}$ Labeled samples of both defective
   and non-defective instances
   are scarce.
$\bm{*}$ Obvious defect features.
$\bm{*}$ Defects are easily distinguishable
   from the background.
$\bm{*}$ Relatively few types of defects.
$\bm{*}$ Low development cost
   for the detection algorithm.
$\bm{*}$ High efficiency in real-time detection.
$\bm{*}$ Low cost for system setup
   and maintenance.
$\bm{*}$ The model’s performance is generally low.
$\bm{*}$ It cannot provide precise defect
   localization and classification.
$\bm{*}$ The model’s robustness and
   generality are poor.
Supervised
Learning
$\bm{*}$ Suitable for scenarios with ample
   annotated training data, including
   samples of both defects and non-defects.
$\bm{*}$ There are not too many restrictions on
   defect types and defect patterns.
$\bm{*}$ High detection accuracy.
$\bm{*}$ Precise defect localization
   and classification.
$\bm{*}$ The model exhibits good robustness.
$\bm{*}$ Dependent on a large amount of
   labeled data.
$\bm{*}$ Difficult to adapt to new types of defects,
   requiring extensive re-labeling of data.
$\bm{*}$ Long development cycle for
   the detection model.
$\bm{*}$ High system setup and maintenance costs.
Semi-supervised
Learning
Only a small fraction of the dataset
has labels, while the majority
of samples are unlabeled.
$\bm{*}$ Utilizes unlabeled data to
   improve performance
   in situations where labeled data is scarce.
$\bm{*}$ Requires fewer labeled samples
   compared to supervised learning.
$\bm{*}$ Utilization of unlabeled data
   may not be fully effective,
   and performance remains constrained
   by the availability of labeled data.
$\bm{*}$ Quality of unlabeled data can
   impact model performance.
Reinforcement
Learning
$\bm{*}$ Applicable to scenarios that
   require autonomous learning of
   defect features.
$\bm{*}$ The defect detection system is capable of
   autonomously or semi-autonomously
   interacting with the inspection
   environment for learning.
$\bm{*}$ Adapts to dynamic environments,
   offering advantages in real-time
   defect detection.
$\bm{*}$ Allows continuous improvement through
   interaction with the environment.
$\bm{*}$ Requires extended training time,
   with high computational resource demands.
$\bm{*}$ Initial instability may be observed,
   necessitating prolonged training for stability.
Generative
Learning
Applicable in situations where there are
relatively few defect samples or
styles of defect types, aiming to
increase the quantity of defect samples or
introduce variations in defect patterns.
$\bm{*}$ Used for synthesizing defect data,
   enhancing the diversity of training data.
$\bm{*}$ Aids the model in learning defect
   features, improving robustness.
$\bm{*}$ Training generative models itself may
   demand significant computational resources.
$\bm{*}$ The quality of generated data significantly
   influences the final model performance,
   requiring careful tuning of
   the generative model.
In material defect detection, the use of different learning methodologies is crucial. Each methodology is designed for specific scenarios and has its own advantages and disadvantages. These are summarized in Table 4, which provides a comparative analysis of the different approaches.
Unsupervised learning is a practical choice when there is a shortage of labelled training data. This approach relies on the algorithm’s ability to identify patterns and anomalies in the data without explicitly labelled examples. The main advantage of unsupervised learning is its flexibility in handling real-world scenarios where obtaining labels for defective and non-defective samples is difficult. This may be due to high costs, time constraints, or the impracticality of manual labelling in complex or nuanced cases. However, this approach is not without limitations. The absence of labelled data, while offering a degree of flexibility, also introduces a potential shortfall in performance metrics. Without explicit labels to guide the learning process, unsupervised algorithms may struggle to achieve the level of precision and accuracy that is attainable with supervised methods. The performance of unsupervised learning algorithms is inherently contingent upon their ability to infer the underlying structure of the data and distinguish between normal variations and genuine defects. This task can be particularly daunting in the context of material defect detection, where the distinction between a defect and a non-defect can be subtle and highly context-dependent.
Supervised learning plays an important role in material defect detection, especially in situations where abundant labelled data is available. It is highly regarded for its effectiveness in achieving high levels of detection accuracy, as well as precise defect localization and classification. The key to its proficiency lies in its utilization of labelled datasets, where each example is annotated with the correct output. This enables the algorithm to learn the complex relationships between input features and desired output labels. One of the strengths of supervised learning is its robustness and predictability. By training on a comprehensive dataset that adequately represents the characteristics of both defects and non-defects, supervised learning models can develop a nuanced understanding of the features that differentiate defects from normal variations in materials. This enables precise defect identification, including size, shape, and location. In turn, this not only facilitates defect detection but also provides valuable insights into their nature and severity. However, the effectiveness of supervised learning is contingent upon the availability and quality of labelled data. Obtaining such data can be labour-intensive and costly, requiring expert knowledge to accurately identify and label defects across potentially large datasets. This is particularly challenging in industrial contexts, where defects can vary greatly and the materials themselves can present a complex range of features that must be correctly interpreted. Furthermore, the dynamic nature of industrial environments and the continuous evolution of materials and manufacturing processes often necessitate frequent updates to the training datasets in order to maintain the accuracy and relevance of supervised learning models. The ongoing need for fresh and accurately labelled data can pose significant logistical and financial challenges, potentially limiting the scalability and adaptability of supervised learning approaches in rapidly changing or resource-constrained settings.
Semi-supervised learning emerges as a particularly effective strategy in situations where labelled data is scarce or expensive to obtain. This approach adeptly bridges the gap between supervised and unsupervised learning by leveraging a small amount of labelled data in conjunction with a larger pool of unlabeled data. The underlying premise is that the unlabeled data, despite the absence of explicit annotations, contains valuable information about the structure and distribution of the data that can be exploited to improve the learning algorithm’s performance. In the context of material defect detection, semi-supervised learning can significantly enhance the model’s ability to identify and classify defects with fewer labelled examples. This is achieved by utilizing the unlabeled data to better understand the feature space’s overall structure, aiding in generalization from the limited labelled samples. The approach can lead to more robust models that are capable of handling the variability and complexity often encountered in material surfaces, without the need for extensive annotated datasets.
Reinforcement learning, on the other hand, presents a paradigm fundamentally different from the traditional supervised or unsupervised learning frameworks. It is particularly suited to interactive and dynamic environments where an agent learns to make decisions by performing actions and receiving feedback in the form of rewards or penalties. This feedback loop enables the agent to learn from its interactions with the environment and optimize its behaviour towards achieving specific goals. In autonomous robotic inspection for defect detection, reinforcement learning can be invaluable. It allows the system to improve its inspection strategies based on the outcomes of its actions, potentially leading to more efficient and effective defect detection processes. However, the training process can be considerably time-consuming and resource-intensive as it requires the agent to explore a vast space of possible actions and their consequences. This exploration process makes it less feasible in scenarios where rapid deployment or real-time decision-making is critical.
Generative learning models, particularly those based on generative adversarial networks (GANs) and variational autoencoders (VAEs), excel in synthesizing realistic defect data, which can significantly enhance the robustness and generalizability of learning algorithms. By generating synthetic data that mimics real-world defects, generative models can augment existing datasets, especially in scenarios where the collection of extensive real defect samples is impractical or impossible due to the rarity or the destructive nature of the defects. However, the deployment of generative learning models is not without its challenges. The process of training these models, particularly GANs, involves a delicate balance between the generator and discriminator components, which requires careful tuning of the model architecture and training parameters. Moreover, generative models are often computationally intensive, requiring significant resources for training and optimization. This can pose a barrier to their application, particularly in resource-constrained environments or scenarios requiring rapid model development and deployment.
4. Techniques for Composite Material Defect Detection
Composite materials are widely used in high-tech fields such as aerospace, military, etc., Due to its advantages such as lightweight, high strength, and corrosion resistance. However, compared to traditional metal materials, composite materials are more prone to various types of defects, as shown in Table 5, which can lead to a decrease in material performance or even failure. Therefore, it is crucial to conduct defect detection on composite materials.
Machine learning techniques for defect detection in composite materials differ significantly from those used for other types of materials. Composite materials require models capable of handling heterogeneous and complex data due to varied material properties and defect types, requiring advanced feature extraction and more complex models, such as deep learning architectures. Multi-modal data integration is often essential, combining information from various sensors, whereas other materials typically rely on single-mode data. The need for diverse and extensive training datasets is more pronounced in composite materials, which often encounter imbalanced datasets and demand techniques like synthetic data generation. Algorithms for composite materials must be highly robust and adaptable to variations in defect presentation and environmental conditions, and capable of real-time processing. Additionally, the complexity of these models poses challenges for interpretability, increasing the need for techniques to explain model decisions. Transfer learning and domain adaptation are also more crucial for composite materials to apply knowledge from other contexts, whereas simpler models and standard training techniques are often sufficient for other materials.
Non-destructive testing (NDT) technology is one of the key methods for evaluating defects in composite materials. Due to the heterogeneity and anisotropy of composite materials, traditional metal material testing techniques are not always applicable. Common defect detection techniques for composite materials include ultrasonic testing, X-ray testing, thermographic imaging, electronic speckle pattern interferometry, acoustic emission testing, and optical non-destructive testing. The following discussion will delve deeper into these methods.
Ultrasonic Testing is commonly used as one of the significant methods to realize composite defect detection. (mcknight20243, ) develops a deep learning approach with 3-D convolutional neural networks (CNNs) for detecting defects in carbon fibre-reinforced polymer composites using volumetric ultrasonic data. To manage the high costs of experimental data, the study uses synthetically generated volumetric data, enhancing the model’s ability to utilize spatial and temporal information. Three architectures were tested: a standard model, a custom model with cuboidal kernels suited for ultrasonic data, and an optimized model developed through neural architecture search (NAS). The NAS-derived model performed best, achieving a significant 22.4% improvement in accuracy over other models and processing data in under 0.5 seconds, proving highly efficient for real-time defect detection. Then, (lu2024research, ) discusses enhancing ultrasonic testing for defect detection in industrial applications, highlighting its current focus on surface scanning and internal defect classification. The study utilizes finite element simulation software to detect simple defects in specimens, with accuracy validated through ultrasonic experiments. Features are extracted using signal processing methods, and a neural network with a Gaussian weight fusion model is used to visualize results. The approach extends to training on signals from simple defects to detect and image complex defects, aiming to improve the accuracy of imaging internal defect morphology. Meanwhile, (oztacs2024image, ) examines the detection of internal damages in composite materials, crucial for industries like aircraft manufacturing. Using glass and carbon fibre-reinforced epoxy composites with artificially induced damages, the materials were subjected to impact tests and analyzed with ultrasonic scanning. The results revealed significant differences in damaged areas, especially in glass fibre-reinforced composites, where the damaged area was up to five times larger than in undamaged samples. Carbon fibre composites with artificial damages showed perforations from impact tests, unlike undamaged samples. Image processing algorithms effectively detected and quantified these barely visible damages, demonstrating the effectiveness of the applied non-destructive testing methods. Besides, X-ray Testing also plays an important role in performing the composite material defect detection. (gong2020deep, ) propose the use of deep transfer learning to enhance the detection of defects in X-ray images of aeronautic composite materials. This approach addresses the limitations of traditional non-destructive testing methods, which are often manual and subjective. By leveraging a model trained on limited samples, the authors achieve 96% classification accuracy, demonstrating the model’s effectiveness in automatic defect detection.
Table 5. Common Composite Material Defects.
Defect Type Defect Manifestation Defect Formation Cause
delamination (de2024understanding, ; liu2020data, )
separation of layers
within the composite material
inadequate bonding during manufacturing
or by mechanical impact
porosity (habibi2021effect, ; al2020effects, )
presence of voids or air pockets
within the material
improper curing, resin infusion, or insufficient
vacuum pressure during manufacturing.
fiber misalignment (zhang2021fibre, )
misalignment or disorientation of reinforcing
fibers within the composite
improper layup or handling
during manufacturing
fiber breakage (suriani2021delamination, ; zhang2021fibre, )
breakage or damage to
individual reinforcing fibers
excessive mechanical stress or impact
matrix cracking (fu2022review, ; mardanshahi2020detection, )
cracks or fractures within the resin
matrix of the composite material
thermal cycling, mechanical loading, or
environmental factors such as moisture ingress
Impact damage (zimmermann2020review, ) surface or subsurface damage impact from foreign objects
inclusions (gong2020deep, ; roy2021compound, )
foreign particles or contaminants
embedded within the composite material
contaminants or impurities introduced during
the melting and solidification processes
fiber pull-out (hemmatian2023machine, ; de2021novel, )
Partial or complete removal of reinforcing
fibers from the matrix
inadequate bonding or insufficient
fiber-matrix adhesion
Another detection technical approach is based on thermographic Imaging for composite material defects. (liu2024multi, ) presents a new vision transformer model for detecting and quantifying defects in composites using thermography. It improves data handling with a novel sampling strategy for infrared data and uses multiple-feature fusion to enhance defect detection. The model, leveraging self-attention mechanisms, excels in accuracy and generalization, effectively identifying defects across composite images. (zhang2024one, ) introduces a novel non-destructive testing method, one-dimensional deep convolutional autoencoder active infrared thermography (1D-DCAE-AIRT), for enhancing defect detection in fibre-reinforced polymer composites. Traditional active infrared thermography methods face challenges with uneven heating, affecting defect visibility. The 1D-DCAE-AIRT method preprocesses thermal images, extracts high-level features using a deep convolutional autoencoder, and uses these features to improve the visualization of defects. Tests on composite specimens with artificially introduced defects showed that this method surpasses traditional techniques like FFT, PCA, ICA, and PLSR in revealing internal defects. (deng2024attention, ) introduces a spatiotemporal-based deep learning model with an attention mechanism designed to analyze thermograms more effectively. The model processes thermal image sequences to characterize barely visible impact damages (BVID) in composite materials, linking these damages to specific impact energy levels. Validated with 100 composite specimens subjected to varying impact forces, the model achieves a classification accuracy of over 95%. This approach offers significant potential for Structural Health Monitoring (SHM), particularly in the aviation industry, by facilitating targeted predictive maintenance and enhancing safety through accurate impact damage analysis.
Optical non-destructive testing has been a promising way to conduct effective and efficient defect detection for composite materials. (palka2024automatic, ) introduces a method using terahertz time-domain spectroscopy (TDS) to automatically detect defects in glass fibre-reinforced polymer (GFRP). Employing a high-speed TDS scanner with an advanced sampling system, the method successfully detected 87% of inclusions and 100% of delaminations in GFRP specimens. A thresholding algorithm based on Gaussian fitting of histograms helped pre-select potential defects, also providing depth information. The results highlight the system’s effectiveness in defect detection in GFRP. (mahmoud2024enhancing, ) proposes a method using hyperspectral imaging (HSI) to detect defects in carbon fibre composites (CFC), which are crucial in various industries for their strength and stiffness. By employing HSI, the method captures detailed spectral data that highlights differences in reflectance between defective and non-defective areas, particularly in the 440 nm to 600 nm range. Integrating HSI with k-means clustering automates defect characterization, enhancing detection speed and accuracy. The approach uses specific wavelengths as markers to measure fracture depths and produces 3D images of defects, offering a precise and efficient tool for evaluating CFC components in industrial applications. (kazemi2024preliminary, ) presents a prototype intelligent quality inspection tool that uses camera vision to monitor fibre distortion in fabric composites during forming processes. Utilizing data from about 30 samples and over 1200 images, the tool employs machine learning algorithms, specifically K-nearest neighbours and Support Vector Machine, to identify defects. These models achieved an error rate of less than 5% in differentiating between defective and non-defective parts, proving effective even amid background noise and image blurriness. This AI tool can be used in a smart factory by operators or robots to scan and assess formed parts efficiently.
More detection techniques include acoustic emission testing and electronic speckle pattern interferometry. (lan2024impact, ) details a method using local defect resonance (LDR) to detect barely visible impact damages (BVIDs) in woven carbon fibre-reinforced polymer (CFRP) laminates through laser-ultrasonic guided waves. The technique involves exciting broadband guided waves with an Nd: YAG pulsed laser, directed using a galvano mirror system, while a fixed broadband piezoelectric sensor records the waves. The recorded data are then processed to visualize the wavefield and extract LDR frequencies, producing high-contrast images of BVIDs. Validated on a CFRP laminate with a single BVID, the method outperformed traditional root mean square (RMS) techniques and successfully demonstrated the detection of multiple BVIDs. (tu2024infrared, ) presents an infrared-induced laser shearography (IILS) method that improves defect detection in composite structures by capturing dynamic deformation data. Utilizing the temperature-deformation connection theory, the method investigates the relationship between dynamic deformations and defects through simulations and experiments. Feature extraction algorithms, particularly tensor robust principal component analysis (TRPCA), are employed to denoise and enhance imaging. This method significantly improves image quality and detection capabilities, validated by metrics such as two-dimensional entropy, contrast, and signal-to-noise ratio. Overall, IILS enhances the reliability and stability of non-destructive testing in industrial applications. (yoon2023defect, ) introduce a non-destructive technique using deep learning to detect delamination in laminated AS4/PEEK composites. The approach utilizes highly nonlinear solitary waves (HNSWs) captured by a granular crystal sensor, which feeds data into an optimized convolutional neural network (CNN). By adjusting CNN parameters and incorporating multiple pre-processed signal modes, the study achieves over 90% accuracy in detecting defects, showcasing the potential for real-time, accurate defect detection in composites.
5. Open Issues and Future Research Directions
Based on a thorough analysis and summary of existing material defect detection technologies, it is evident that although current machine learning algorithms contribute significantly to the improvement of material defect detection performance, there are still notable technological barriers and performance shortcomings when addressing the rigorous demands of real-world industrial applications. These applications necessitate high levels of real-time performance, stability, scalability, and cost-effectiveness. However, existing solutions often struggle to meet these requirements simultaneously, limiting their widespread adoption in industry. In this section, we will outline key challenges and give potential future research directions concerning current material defect detection technologies, which help enable the development of more efficient, accurate, and reliable defect detection systems that meet the rigorous demands of real-world industrial scenarios.
5.1. Establishment of Standardized Datasets for Material Defects
In the field of material defect detection, a significant challenge arises from the lack of standardized and widely accepted datasets. The absence of a standard dataset makes it difficult to achieve consistent evaluations of algorithms used in different studies. This hinders the comparability of research results and impedes the further development of the field. Additionally, the absence of diverse and extensive datasets poses challenges to the generalization ability of developed algorithms in practical applications.
Therefore, it is important to create standardized and diverse material defect detection datasets that encompass various types of materials, multiple defect categories, and images with different scales and resolutions to ensure comprehensive testing of algorithms in various scenarios. One potential solution is to encourage industry collaboration for the joint establishment and maintenance of standardized datasets. By sharing data, research institutions and businesses can better assess their respective algorithms, promoting collaborative development across the entire field. Additionally, advocating for the creation of open-source platforms enables researchers and practitioners to share algorithms, data, and experimental results, thereby advancing the open development of material defect detection technologies.
5.2. Advanced Techniques for Complex Material Defect Detection
Existing defect detection technologies usually focus on a singular type of defect. However, in real-world scenarios, materials often exhibit various types of defects simultaneously, and each type of defect may possess distinct shapes, sizes, and characteristics. Additionally, the interaction of multiple defects can further complicate detection tasks. For example, the presence of one defect may obscure the features of another defect. Furthermore, different types of defects may exhibit an imbalanced distribution in practical applications, potentially resulting in poor model performance on certain categories. This adds complexity and challenges to the task of detecting complex defects but is crucial for meeting the requirements of real-world material defect detection applications.
To address multi-category defect detection, a multi-task learning framework can be employed to simultaneously train models to handle multiple defect categories. This helps the model better understand the relationships between defects and improves overall performance. Furthermore, integrating multiple sensors or data sources, such as optical images, thermal imaging, acoustic waves, etc., can provide more comprehensive and multi-angle information, which helps enhance the ability to handle complex defects.
5.3. Small Sample Material Defect Detection
In real-world applications of material defect detection, the advancement of manufacturing processes has led to a scarcity of defect samples compared to normal samples. Moreover, annotating defect samples requires a significant amount of human effort, further exacerbating the rarity of defect samples. This scenario is referred to as small sample material defect detection. Small sample material defect detection involves effectively detecting defects when there is an insufficient quantity of samples in the training data. However, most existing deep learning methods fall under the category of large sample learning and typically perform poorly in small sample situations. This results in models that struggle to accurately generalize to unseen defect samples, leading to low detection rates and unstable model performance, ultimately failing to meet practical inspection requirements.
Small sample defect detection poses a challenging problem due to the limitation of training data. To address this challenge, potential approaches include transfer learning, meta-learning, generative adversarial networks (GANs), active learning, reinforcement learning, and synthetic datasets. Transfer learning enhances a model’s generalization ability in small sample contexts by leveraging knowledge pre-trained on other tasks. Meta-learning trains a model to handle multiple tasks, improving its adaptation to new defect detection tasks. GANs generate synthetic defect samples to augment training data. Active learning reduces annotation requirements by selecting samples with high information value. Reinforcement learning enhances performance through model interaction with the environment. Synthetic datasets create simulated defect data to increase training samples. The choice of these methods depends on the nature of the problem and available resources. Through their application, the performance and adaptability of small sample defect detection models can be improved, better meeting the diverse conditions encountered in actual industrial production.
5.4. Dynamic Adaptation to Evolving Defect Patterns
As manufacturing technology advances, the dynamic emergence of novel defects may pose challenges to existing defect detection methods within the current context. Firstly, the unknown shapes, features, or variations of novel defects make it difficult for existing models to accurately generalize to these changes. Secondly, the lack of standards or specifications for novel defects prevents existing detection methods from effectively adjusting based on prior experience. Additionally, the annotation data for novel defects may be extremely limited, leading to difficulties for existing supervised learning methods in learning and adapting. Finally, addressing novel defects may require more real-time monitoring and detection, a demand that existing methods may not be able to meet.
To address the challenge of inapplicability in the current context, various potential solutions can be considered. Firstly, adopting incremental learning methods enables the model to update continually as it encounters new samples, adapting to the changes presented by novel defects. Secondly, utilizing adaptive algorithms allows the model to automatically adjust parameters based on the appearance of novel defects, enhancing adaptability to unknown defects. Transfer learning is another viable solution, transferring the model’s experience from known defects to the detection of novel defects. Additionally, integrating multiple sensors or data sources, such as images, sounds, temperatures, etc., to form a multimodal fusion system can provide a more comprehensive, multi-angle perception of novel defects. By adopting these solutions, the capability of defect detection methods to dynamically adapt to novel defects can be enhanced, addressing the evolving demands of manufacturing technology.
5.5. Real-time Online Defect Learning and Detection
With the advancement of manufacturing technology, the demand for real-time defect learning and inspection continues to grow. Existing defect detection methods predominantly employ offline learning, wherein the model is trained using a pre-prepared dataset during the training phase and subsequently deployed to the production line. However, this approach faces limitations in the early stages of training due to a lack of annotated defect samples, making it challenging for the model to achieve optimal performance in actual deployment. In practical applications, production lines continually generate new samples and may introduce new types of defects, rendering offline learning methods less flexible. Therefore, employing online learning methods to adapt the model in real-time becomes a key issue in both research and application.
To address these potential issues, a series of solutions can be implemented. Firstly, incorporating incremental learning enables the model to be updated in real-time upon receiving new samples, enhancing real-time adaptability to novel defects. Secondly, optimizing algorithms and hardware infrastructure involves employing efficient computational methods and advanced hardware technologies to meet the computational requirements of real-time defect detection. Additionally, stream data processing techniques can reduce the latency of sensor data reaching the detection system, improving real-time performance. Through the combined application of these solutions, it is possible to effectively address the issues of insufficient annotated data and the recognition of new defect types encountered by offline learning methods in practical applications. This makes the defect detection system more flexible, robust, and capable of real-time operation.
5.6. Composite Material Defect Detection
Composite materials are typically composed of multiple materials, which introduces heterogeneity and makes detection more challenging due to varying acoustical, thermal, and electrical properties among different constituents. Furthermore, composites often exhibit complex geometries and structures, and defects may be located in areas that are difficult for sensors to access, thereby significantly increasing the difficulty of detection. Additionally, certain types of composite materials possess high absorbance and scattering characteristics, introducing noise and interference to optical sensing devices, thus potentially reducing the accuracy of defect detection. Lastly, defects in composite materials are often small and irregularly shaped, posing challenges for both detection and localization.
To address the aforementioned challenges in defect detection of composite materials, future research efforts could focus on the comprehensive utilization of multimodal detection techniques, advanced imaging technologies, optimized sensor designs and signal processing algorithms, as well as machine learning and artificial intelligence approaches. These advanced technologies hold promise for improving the accuracy, efficiency, and reliability of defect detection in composite materials.
Various issues, challenges, and possible future solutions for material defect detection are summarized in Table 6.
Table 6. Summary of Issues, Challenges, and Possible Solutions in Material Defect Detection
Issue Challenges Involved Possible Future Solutions
1. Establishment of Standardized Datasets for Material Defects Difficulty in achieving consistent evaluations of algorithms, hinders comparability of research results, impedes field development Create standardized and diverse datasets, encourage industry collaboration for joint establishment and maintenance of datasets, advocate for open-source platforms
2. Advanced Techniques for Complex Material Defect Detection Materials exhibit various defects simultaneously, interaction of multiple defects complicates detection, imbalanced distribution of defects Multi-task learning framework, integrate multiple sensors or data sources, explore multi-modal defect detection models
3. Small Sample Material Defect Detection Insufficient quantity of defect samples, poor model performance in small sample situations Transfer learning, meta-learning, generative adversarial networks (GANs), active learning, reinforcement learning, synthetic datasets
4. Dynamic Adaptation to Evolving Defect Patterns Unknown shapes, features, or variations of novel defects, lack of standards, limited annotation data, real-time monitoring requirements Incremental learning methods, adaptive algorithms, transfer learning, integrate multiple sensors or data sources
5. Real-time Online Defect Learning and Detection Insufficient annotated defect samples, introduction of new defect types, lack of flexibility in offline learning methods Incremental learning, optimizing algorithms and hardware infrastructure, stream data processing techniques
6. Composite Material Defect Detection Heterogeneity in composite materials, complex geometries and structures, noise and interference in optical sensing, small and irregularly shaped defects Multimodal detection techniques, advanced imaging technologies, optimized sensor designs, signal processing algorithms, machine learning and AI approaches
6. Conclusion
In conclusion, this survey paper has demonstrated the increasingly vital role of machine learning techniques in the field of modern materials defect detection. The integration of machine learning with machine vision technologies has led to significant advancements and practical applications in real-world defect detection scenarios. As technology continues to evolve, it is evident that machine learning-driven defect detection will remain a prominent area of research with far-reaching implications for various industries.
This paper has provided a comprehensive review of the latest research developments in materials defect detection, covering task definitions, types of defects, challenges in detection, the framework of machine vision-based detection systems, and a detailed analysis of commonly used machine learning algorithms, including their respective advantages and disadvantages. By offering a thorough exploration of these aspects, this survey serves as a valuable resource for both researchers and practitioners in the field.
Looking ahead, the landscape of materials defect detection is dynamic and continually evolving, presenting both challenges and opportunities. This paper has identified current research challenges and potential future directions, aiming to bridge the gap between theoretical research and industrial application. By highlighting these areas, we hope to inspire and guide future researchers, contributing to ongoing development and innovation in materials defect detection. Furthermore, this survey provides practical insights that are invaluable for industrial practitioners. The comprehensive analysis of machine learning algorithms and their applications in defect detection can inform decision-making and strategic planning, enabling organizations to leverage the power of these technologies effectively. By understanding the capabilities and limitations of various approaches, practitioners can make informed choices and develop robust defect detection systems tailored to their specific needs.
Acknowledgements.
This work is supported by the Centre for Future Materials at the University of Southern Queensland, Australia and Air Force Research Laboratory (AFRL), USA.